# TWISTED CONJUGACY IN DIHEDRAL ARTIN GROUPS I: TORUS KNOT GROUPS

### v3

@@ INTRODUCTION
for some $1\leq i\leq\tau$. We consider each of the 4 conditions from 3.8, when applied to $u^{\prime}$.
@@
$\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)=\{u\in\left(3^{++}\right)\cup\left(3^{--}\right)\cup\left(3^{0*}\right)\mid u\;\text{starts and ends with opposite factors}\}.$
@@
Proof.
By the previous discussion, we only need to consider words $u\in\overline{(3)}$ which contain any $y^{\pm(k+1)}$ terms. For $\epsilon=\pm 1$, we can first rewrite all pairs $\left(y^{\varepsilon(k+1)},y^{-\epsilon(k+1)}\right)$ as $\left(y^{-\epsilon k},y^{\epsilon k}\right)$, until we are left only with $y^{k+1}$ terms or $y^{-(k+1)}$ terms (see A.1). Then, we can take a $\phi$-cyclic permutation, which switches all $y^{\epsilon(k+1)}$ terms to $y^{-\epsilon(k+1)}$, and then apply the following rewrite rule to obtain a word $u^{\prime}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$:
@@
where $\epsilon=\pm 1,s,t,z\in X^{\ast}$.
Proof.
For any $y^{b}$ factor in $u$, we have that $|b|\leq k$ by 3.9, and so Condition 2 of 3.11 is never satisfied. For the first condition, the only rewrite rule we can apply is Eq. 5, for all pairs of $\left(x,x^{-1}\right)$ terms. ∎
Remark 3.13.
@@
Proof.
If $u$ is uniquely geodesic we are done, so suppose $u$ is a non-unique geodesic. From 3.12, we only need to consider rewriting of the form in Eq. 5. Assume $u$ is of the form $u=sxtx^{-1}z$, for some $s,t,z\in X^{\ast}$, and $u^{\prime}=sx^{-1}txz=_{G}u$. We now check if all $\phi$-cyclic permutations of $u^{\prime}$ are geodesic. All $\phi$-cyclic permutations of $u^{\prime}$ are equivalent to a $\phi$-cyclic permutation of $u$, which are geodesic by assumption, except for the following cases.
@@
A symmetric argument holds in the case where $u=sx^{-1}txz$. ∎
@@
Proof.
The reverse implication is clear by 2.11 and Lemma 3.14. For the forward direction, assume $v=_{G}\phi(w)^{-1}uw=\mathsf{rev}(w)uw$ for some $w\in X^{\ast}$. We assume $w$ is written in geodesic form, as defined in Table 2. We consider different cases for $w\in X^{\ast}$.
@@
for some $b_{i}$ where $-k\leq b_{i}<0$, and some $s,t\in X^{\ast}$. Since $b_{i}+m\geq k+1$, and no $y^{\beta}$ term, where $|\beta|>k$, can occur in $v$ by 3.9, then either this type of cancellation cannot occur, or we need to reverse this rewrite. Since $c>0$, the only way we can do this is through the existence of an $x^{-1}$ term in $u$. If this is the case, we can apply the following rewrite:
@@
Let $u=_{G}u_{1}u_{2}$ such that $u_{1},u_{2}\in X^{\ast}$ are geodesics, such that all $x$ factors with exponents in $I_{1}$ lie in $u_{1}$, and similarly $x$ factors with exponents in $I_{2}$ lie in $u_{2}$. By applying $x^{-1}\Delta=x$ to all $x$ factors in $u$ with exponents in $I$, we can write $v$ in the form $v=_{G}u_{1}u_{2}\Delta^{2c}=_{G}v_{1}v_{2}$, for some geodesic words $v_{1},v_{2}\in X^{\ast}$ such that $v_{1}v_{2}\in X^{\ast}$ is also geodesic. Note $u_{1}u_{2},v_{1}v_{2}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 3.14. We now have
@@
Again note $\overline{\phi(v_{2})}\overline{v_{1}}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 3.14. Finally, we can apply another $\phi$-cyclic permutation to obtain a word equivalent to $u$:
@@
Suppose $v=_{G}\mathsf{rev}(w)uw$, where $w$ is Garside free. Since $v$ starts and ends with opposite factors, there exists free cancellation in $\mathsf{rev}(w)uw$. Moreover, this cancellation occurs either in $\mathsf{rev}(w)u$ or $uw$, but not both, since $u$ starts and ends with opposite factors. Assume free cancellation occurs in $uw$. We let $u=_{G}u_{1}u_{2},\;w=_{G}u^{-1}_{2}z.$ Then we can write $v=_{G}\mathsf{rev}(z)\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}z$. Since $u_{1}u_{2}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 3.14, the subword $\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}=\phi(u_{2})u_{1}$ must be geodesic. Assuming $u_{2}$ is maximal, we can assume $z$ is empty, since $v$ starts and ends with opposite factors. Therefore $v=_{G}\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}$ and we obtain our required sequence of $\phi$-cyclic permutations and equivalent geodesics as follows:
@@
after $2c$ $x^{-1}$ terms from $\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}$ have cancelled with the Garside element. Since $v$ starts and ends with opposite letters, we can assume $v=_{G}\overline{\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}}$, and from here we can apply the same proof as Case 1. A similar proof holds when free cancellation occurs in $\mathsf{rev}(w_{1})u$, or when we rewrite factors of $\mathsf{rev}(w_{1})uw_{1}$. ∎
Proposition 3.16.
@@
    $\phi\colon x\mapsto x^{-1},y\mapsto y^{-1}\in\mathrm{Out}(G(m))$.
Step 1: Geodesic form {adjustwidth}1.5cm Write $u,v$ in geodesic forms $\overline{u},\overline{v}$ respectively, given by Table 2. Step 2: Simple cases {adjustwidth}1.5cm If $\overline{u}$ or $\overline{v}$ is of the form $x^{a}$ or $y^{b}$, then apply Lemma 3.3 and 3.15 to determine if $\overline{u}$ and $\overline{v}$ are twisted conjugate. In particular, $\overline{u}\sim_{\phi}\overline{v}$ if and only if:
@@
Otherwise, we can assume $\overline{u},\overline{v}$ contain at least one $x$ and $y$ factor. Step 3: Twisted cyclic geodesics {adjustwidth}1.5cm Apply $\phi$-cyclic permutations and geodesic reductions to obtain $\overline{u}_{\min},\overline{v}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $\overline{u}\sim_{\phi}\overline{u}_{\min}$ and $\overline{v}\sim_{\phi}\overline{v}_{\min}$ (existence follows from 3.2 and Lemma 3.10). If $l(\overline{u}_{\min})\neq l(\overline{v}_{\min})$, then Output = False (by 3.15). Step 4: Set of representatives {adjustwidth}1.5cm Compute all possible sequences of $\phi$-cyclic permutations and equivalent words from $\overline{u}_{\min}$, to get a finite set $\mathcal{D}$ of minimal length twisted conjugacy representatives. If $\overline{v}_{\min}\in\mathcal{D}$, then Output = True. Otherwise, Output = False (by 3.15). ∎
Example 3.17.
@@
$G(3)=\langle x,y\>|\;x^{2}=y^{3}\rangle$, and the word $u=x^{-1}yx^{-1}y\in X^{\ast}$. Here $u\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by 3.9, and Fig. 1 gives the finite set $\mathcal{D}$ of words obtained by computing all possible $\phi$-cyclic permutations and equivalent words. Here the arrowed lines indicate $\phi$-cyclic permutations.
@@
Figure 1: Set $\mathcal{D}$ obtained in 3.17.
Complexity of algorithm
We make some observations to improve our algorithm, which will lead to linear time complexity. We recall 3.17 as a motivating example. Rather than computing all words of minimal length in $\mathcal{D}$, we could have instead computed all $\phi$-cyclic permutations from $u\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ only. Then, if we want to check if $v\in\mathcal{D}$, we simply check if $v$ is equal, as group elements, to a word $u^{\prime}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained from $u$ via $\phi$-cyclic permutations. This step runs in linear time, based on the length of $u$, and we will go on to show that this idea can be applied in all cases.
@@
Proof.
By 3.15, $u\sim_{\phi}v$ if and only if $u$ and $v$ are related by a sequence of $\phi$-cyclic permutations and equivalent geodesics. Since Eq. 5 does not affect any $y$ factors, we can assume $y$ factors are changed only by $\phi$-cyclic permutations. ∎
3.18 allows us to add the following information to Step 3 of our algorithm.
@@
Proof.
First recall from 3.13 that to check if two words $u,v\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ represent the same group element, then we only need to check that the exponents of $y$ factors of $u$ and $v$ match exactly, and that the number of $x$ terms in $u$ equals the number of $x$ terms in $v$ (or analogously, the number of $x^{-1}$ terms in $u$ equals the number of $x^{-1}$ terms in $v$). Combined with 3.15, we can find unique representatives for each group element in a twisted conjugacy class as follows.
@@
In total, we obtain $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements, by computing all $\phi$-cyclic permutations of $\overline{u}_{\min}$ using this block decomposition. To see this, we group together block decomposition’s which only differ by one $\phi$-cyclic permutation of a block. These pairs have either odd or even number of $x$ terms, which gives us $m+1$ group elements. Summing across all $y$ powers, to include all possible $\phi$-cyclic permutations, we get $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements. Since $m<n$, this value has an upper bound of $n(n+1)$. ∎
@@
if $\tau$ is even, and define analogously for $m$ odd. We compute all $\phi$-cyclic permutations of $\overline{u}_{\min}\mathsf{CycGeo}_{\phi,\overline{(3)}}$ using the block decomposition, whilst keeping track of how many $x$ terms can exist for each element we obtain (for each $\phi$-cyclic permutation, we switch between odd and even). For each group element $w\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained via $\phi$-cyclic permutations from $\overline{u}_{\min}$, we check whether the exponents of $y$ factors of $w$ and $\overline{v}_{\min}$ match exactly, and that $X(w)=X(\overline{v}_{\min})$. If this holds for some $w$, then Output = True. Otherwise, if no such $w$ exists, then Output = False .
Example 3.20.
Recall 3.17. Here $m=2$ and $(m+1)(m+\sum^{m}_{i=1}|b_{i}|)=3(2+2)=12$ from 3.19, and so there are 12 minimal length group elements within the set $\mathcal{D}$. We list them in Table 1, by taking the minimal shortlex representative for each group element (where $x<x^{-1}<y<y^{-1})$, and also recording the parity of the number of $x$-terms in each group element.
@@
Table 1: Group elements from 3.17
@@
Consider $w=yxy^{-1}x$ from this list. Then the $y$-exponents of $w$ and $\overline{v}_{\min}$ match exactly, and $X(w)=X(\overline{v}_{\min})=\{0,2\}$, and so our algorithm will output True as required. Alternatively, consider $\overline{v}_{\min}=yxy^{-1}x^{-1}$. Again by Fig. 1, $\overline{u}_{\min}\not\sim_{\phi}\overline{v}_{\min}$. Here $\{0,2\}=X(w)\neq X(\overline{v}_{\min})=\{1\}$, and since $w$ is the only word from this list with matching $y$-exponents to $\overline{v}_{\min}$, then our algorithm will output False as required.
We are now able to determine linear complexity for our adapted algorithm.
Proposition 3.21.
@@
Proof.
Let $u,v\in X^{\ast}$ be our input words, where $|u|+|v|=n$. Step 1 of our algorithm runs in linear time, by checking the conditions from Table 2, and applying the relations from $G(m)$ (see A.1). Linearity is also clear in Step 2.
@@
Next we need to understand the complexity of finding a minimal length element in $\mathsf{CycGeo}_{\phi,\overline{(3)}}$, using Lemma 3.10. This only requires us to check for any $y^{\pm(k+1)}$ terms, and apply $\phi$-cyclic permutations and rewriting rules. The complexity of this step is bounded above by $2n$.
@@
First, we adjust the input tape, so that we can determine if the $y$-terms are a $\phi$-cyclic permutation of each other. Then we add a counter which, when reading $x^{\pm 1}$ terms, keeps track of when they differ from the symbol on the pushdown list. The 2DPDA accepts a word if and only if the $y$-terms are a $\phi$-cyclic permutation of each other, and the counter is equal to 0(mod $0\ (\mathrm{mod}\ 2)$ after reading a word. This extra step of adding a counter to our 2DPDA can be achieved in constant time, and so Step 4 takes linear time overall on a RAM machine. ∎
ALGORITHM FOR THE TWISTED CONJUGACY PROBLEM

### v5

@@ 3.2. Twisted cyclic geodesics
for some $1\leq i\leq\tau$. We consider each of the 4 conditions from Proposition 9, when applied to $u^{\prime}$.
@@
$\mathsf{CycGeo}_{\phi,\overline{(3)}}(G,X)=\{u\in\left(3^{++}\right)\cup\left(3^{--}\right)\cup\left(3^{0*}\right)\mid u\;\text{starts and ends with opposite letters}\}.$
@@
Proof 3.4.
By the previous discussion, we only need to consider words $u\in\overline{(3)}$ which contain any $y^{\pm(k+1)}$ terms. For $\epsilon\in\{\pm 1\}$, we can first rewrite all pairs $\left(y^{\varepsilon(k+1)},y^{-\epsilon(k+1)}\right)$ as $\left(y^{-\epsilon k},y^{\epsilon k}\right)$, until we are left only with $y^{k+1}$ terms or $y^{-(k+1)}$ terms (see Proposition 24). Then, we can take a $\phi$-cyclic permutation, which switches all $y^{\epsilon(k+1)}$ terms to $y^{-\epsilon(k+1)}$, and then apply the following rewrite rule to obtain a word $u^{\prime}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$:
@@
where $\epsilon\in\{\pm 1\},s,t,z\in X^{\ast}$.
Proof 3.5.
For any $y^{b}$ factor in $u$, we have that $|b|\leq k$ by Corollary 10, and so Condition 2 of Proposition 12 is never satisfied. For the first condition, the only rewrite rule we can apply is Eq. 5, for all pairs of $\left(x,x^{-1}\right)$ terms.
Remark 14.
@@
Proof 3.6.
If $u$ is uniquely geodesic we are done, so suppose $u$ is a non-unique geodesic. From Corollary 13, we only need to consider rewriting of the form in Eq. 5. Assume $u$ is of the form $u=sxtx^{-1}z$, for some $s,t,z\in X^{\ast}$, and $u^{\prime}=sx^{-1}txz=_{G}u$. We now check if all $\phi$-cyclic permutations of $u^{\prime}$ are geodesic. All $\phi$-cyclic permutations of $u^{\prime}$ are equivalent to a $\phi$-cyclic permutation of $u$, which are geodesic by assumption, except for the following cases.
@@
A symmetric argument holds in the case where $u=sx^{-1}txz$.
@@
Proof 3.7.
The reverse implication is clear by Definition 2.2 and Lemma 15. For the forward direction, assume $v=_{G}\phi(w)^{-1}uw=\mathsf{rev}(w)uw$ for some $w\in X^{\ast}$. We assume $w$ is written in geodesic form, as defined in Table 2. We consider different cases for $w\in X^{\ast}$.
@@
for some $b_{i}$ where $-k\leq b_{i}<0$, and some $s,t\in X^{\ast}$. Since $b_{i}+m\geq k+1$, and no $y^{\beta}$ term, where $|\beta|>k$, can occur in $v$ by Corollary 10, then either this type of cancellation cannot occur, or we need to reverse this rewrite. Since $c>0$, the only way we can do this is through the existence of an $x^{-1}$ term in $u$. If this is the case, we can apply the following rewrite:
@@
Let $u=_{G}u_{1}u_{2}$ such that $u_{1},u_{2}\in X^{\ast}$ are geodesics, such that all $x$ factors with exponents in $I_{1}$ lie in $u_{1}$, and similarly $x$ factors with exponents in $I_{2}$ lie in $u_{2}$. By applying $x^{-1}\Delta=x$ to all $x$ factors in $u$ with exponents in $I$, we can write $v$ in the form $v=_{G}u_{1}u_{2}\Delta^{2c}=_{G}v_{1}v_{2}$, for some geodesic words $v_{1},v_{2}\in X^{\ast}$ such that $v_{1}v_{2}\in X^{\ast}$ is also geodesic. Note $u_{1}u_{2},v_{1}v_{2}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 15. We now have
@@
Again note $\overline{\phi(v_{2})}\overline{v_{1}}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 15. Finally, we can apply another $\phi$-cyclic permutation to obtain a word equivalent to $u$:
@@
Suppose $v=_{G}\mathsf{rev}(w)uw$, where $w$ is Garside free. Since $v$ starts and ends with opposite factors, there exists free cancellation in $\mathsf{rev}(w)uw$. Moreover, this cancellation occurs either in $\mathsf{rev}(w)u$ or $uw$, but not both, since $u$ starts and ends with opposite factors. Assume free cancellation occurs in $uw$. We let $u=_{G}u_{1}u_{2},\;w=_{G}u^{-1}_{2}z.$ Then we can write $v=_{G}\mathsf{rev}(z)\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}z$. Since $u_{1}u_{2}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Lemma 15, the subword $\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}=\phi(u_{2})u_{1}$ must be geodesic. Assuming $u_{2}$ is maximal, we can assume $z$ is empty, since $v$ starts and ends with opposite factors. Therefore $v=_{G}\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}$ and we obtain our required sequence of $\phi$-cyclic permutations and equivalent geodesics as follows:
@@
after $2c$ $x^{-1}$ terms from $\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}$ have cancelled with the Garside element. Since $v$ starts and ends with opposite letters, we can assume $v=_{G}\overline{\mathsf{rev}\left(u^{-1}_{2}\right)u_{1}}$, and from here we can apply the same proof as Case 1. A similar proof holds when free cancellation occurs in $\mathsf{rev}(w_{1})u$, or when we rewrite factors of $\mathsf{rev}(w_{1})uw_{1}$.
Proposition 17.
@@
    $\phi\colon x\mapsto x^{-1},y\mapsto y^{-1}\in\mathrm{Aut}(G(m))$.
Step 1: Geodesic form {adjustwidth}1.5cm Write $u,v$ in geodesic forms $\overline{u},\overline{v}$ respectively, given by Table 2. Step 2: Simple cases {adjustwidth}1.5cm If $\overline{u}$ or $\overline{v}$ is of the form $x^{a}$ or $y^{b}$, then apply Lemma 7 and Theorem 16 to determine if $\overline{u}$ and $\overline{v}$ are twisted conjugate. In particular, $\overline{u}\sim_{\phi}\overline{v}$ if and only if:
@@
Otherwise, we can assume $\overline{u},\overline{v}$ contain at least one $x$ and $y$ factor. Step 3: Twisted cyclic geodesics {adjustwidth}1.5cm Apply $\phi$-cyclic permutations and geodesic reductions to obtain $\overline{u}_{\min},\overline{v}_{\min}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$, such that $\overline{u}\sim_{\phi}\overline{u}_{\min}$ and $\overline{v}\sim_{\phi}\overline{v}_{\min}$ (existence follows from Proposition 6 and Lemma 11). If $l(\overline{u}_{\min})\neq l(\overline{v}_{\min})$, then Output = False (by Theorem 16). Step 4: Set of representatives {adjustwidth}1.5cm Compute all possible sequences of $\phi$-cyclic permutations and equivalent words from $\overline{u}_{\min}$, to get a finite set $\mathcal{D}$ of minimal length twisted conjugacy representatives. If $\overline{v}_{\min}\in\mathcal{D}$, then Output = True. Otherwise, Output = False (by Theorem 16).
{exa}
@@
$G(3)=\langle x,y\>|\;x^{2}=y^{3}\rangle$, and the word $u=x^{-1}yx^{-1}y\in X^{\ast}$. Here $u\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ by Corollary 10, and Fig. 1 gives the finite set $\mathcal{D}$ of words obtained by computing all possible $\phi$-cyclic permutations and equivalent words. Here the arrowed lines indicate $\phi$-cyclic permutations.
@@
Figure 1. Set $\mathcal{D}$ obtained in Example 3.2.
3.3. Complexity of algorithm
We make some observations to improve our algorithm, which will lead to linear time complexity. We recall Example 3.2 as a motivating example. Rather than computing all words of minimal length in $\mathcal{D}$, we could have instead computed all $\phi$-cyclic permutations from $u\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ only. Then, if we want to check if $v\in\mathcal{D}$, we simply check if $v$ is equal, as group elements, to a word $u^{\prime}\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained from $u$ via $\phi$-cyclic permutations. This step runs in linear time, based on the length of $u$, and we will go on to show that this idea can be applied in all cases.
@@ 3.3. Complexity of algorithm
Proof 3.9.
By Theorem 16, $u\sim_{\phi}v$ if and only if $u$ and $v$ are related by a sequence of $\phi$-cyclic permutations and equivalent geodesics. Since Eq. 5 does not affect any $y$ factors, we can assume $y$ factors are changed only by $\phi$-cyclic permutations.
Proposition 18 allows us to add the following information to Step 3 of our algorithm.
@@
Proof 3.10.
First recall from Remark 14 that to check if two words $u,v\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ represent the same group element, then we only need to check that the exponents of $y$ factors of $u$ and $v$ match exactly, and that the number of $x$ terms in $u$ equals the number of $x$ terms in $v$ (or analogously, the number of $x^{-1}$ terms in $u$ equals the number of $x^{-1}$ terms in $v$). Combined with Theorem 16, we can find unique representatives for each group element in a twisted conjugacy class as follows.
@@
In total, we obtain $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements, by computing all $\phi$-cyclic permutations of $\overline{u}_{\min}$ using this block decomposition. To see this, we group together block decomposition’s which only differ by one $\phi$-cyclic permutation of a block. These pairs have either odd or even number of $x$ terms, which gives us $m+1$ group elements. Summing across all $y$ powers, to include all possible $\phi$-cyclic permutations, we get $(m+1)(m+\sum^{m}_{i}|b_{i}|)$ group elements. Since $m<n$, this value has an upper bound of $n(n+1)$.
@@
if $\tau$ is even, and define analogously for $m$ odd. We compute all $\phi$-cyclic permutations of $\overline{u}_{\min}\mathsf{CycGeo}_{\phi,\overline{(3)}}$ using the block decomposition, whilst keeping track of how many $x$ terms can exist for each element we obtain (for each $\phi$-cyclic permutation, we switch between odd and even). For each group element $w\in\mathsf{CycGeo}_{\phi,\overline{(3)}}$ obtained via $\phi$-cyclic permutations from $\overline{u}_{\min}$, we check whether the exponents of $y$ factors of $w$ and $\overline{v}_{\min}$ match exactly, and that $X(w)=X(\overline{v}_{\min})$. If this holds for some $w$, then Output = True. Otherwise, if no such $w$ exists, then Output = False . {exa} Recall Example 3.2. Here $m=2$ and $(m+1)(m+\sum^{m}_{i=1}|b_{i}|)=3(2+2)=12$ from Proposition 19, and so there are 12 minimal length group elements within the set $\mathcal{D}$. We list them in Table 1, by taking the minimal shortlex representative for each group element (where $x<x^{-1}<y<y^{-1})$, and also recording the parity of the number of $x$-terms in each group element.
@@
Table 1. Group elements from Example 3.2
@@
Consider $w=yxy^{-1}x$ from this list. Then the $y$-exponents of $w$ and $\overline{v}_{\min}$ match exactly, and $X(w)=X(\overline{v}_{\min})=\{0,2\}$, and so our algorithm will output True as required. Alternatively, consider $\overline{v}_{\min}=yxy^{-1}x^{-1}$. Again by Fig. 1, $\overline{u}_{\min}\not\sim_{\phi}\overline{v}_{\min}$. Here $\{0,2\}=X(w)\neq X(\overline{v}_{\min})=\{1\}$, and since $w$ is the only word from this list with matching $y$-exponents to $\overline{v}_{\min}$, then our algorithm will output False as required. We are now able to determine linear complexity for our adapted algorithm.
Proposition 20.
@@
Proof 3.11.
Let $u,v\in X^{\ast}$ be our input words, where $|u|+|v|=n$. Step 1 of our algorithm runs in linear time, by checking the conditions from Table 2, and applying the relations from $G(m)$ (see Proposition 24). Linearity is also clear in Step 2.
@@
Next we need to understand the complexity of finding a minimal length element in $\mathsf{CycGeo}_{\phi,\overline{(3)}}$, using Lemma 11. This only requires us to check for any $y^{\pm(k+1)}$ terms, and apply $\phi$-cyclic permutations and rewriting rules. The complexity of this step is bounded above by $2n$.
@@
First, we adjust the input tape, so that we can determine if the $y$-terms are a $\phi$-cyclic permutation of each other. Then we add a counter which, when reading $x^{\pm 1}$ terms, keeps track of when they differ from the symbol on the pushdown list. The 2DPDA accepts a word if and only if the $y$-terms are a $\phi$-cyclic permutation of each other, and the counter is equal to 0(mod $0\ (\mathrm{mod}\ 2)$ after reading a word. This extra step of adding a counter to our 2DPDA can be achieved in constant time, and so Step 4 takes linear time overall on a RAM machine.
4. Algorithm for the twisted conjugacy problem